# Supporting Creative Ownership through Deep Learning-Based Music Variation

### v1

@@ 1 Introduction
The music industry is currently experiencing an AI-driven transformation, with various companies now providing complete song generation through prompts Nayar (2025). While the revolutionary potential of these technologies remains unclear, a more promising direction has emerged: using AI to support rather than replace human creativity O’Toole and Ágnes Horvát (2024); Krol et al. (2025). These supportive applications, known as Artificial Intelligence Creativity Support Tools (AI-CSTs), assist in the creative process, offering varying degrees of agency and control. Some tools are designed to automate routine or technically challenging tasks such as audio-to-midi transcription Hawthorne et al. (2021) and stem splitting Hennequin et al. (2020). These systems support the creative process but rarely provide creative ideas. Other tools instead offer more creative involvement, assisting in ideation and even the extension of musical ideas. Notable systems include the Multi-track Music Machine (MMM) Ens and Pasquier (2020) which generates controllable multi-track MIDI; Music Transformer Huang et al. (2019), offering melody harmonisation; and the Anticipatory Transformer Thickstun et al. (2023), which supports infilling of musical phrases.
@@
Although some research has demonstrated the potential for AI-CSTs to support users Louie et al. (2020), recent work has highlighted the potential pitfalls of using AI for ideation. For example, Wadinambiarachchi et al. (2024) demonstrated how AI promotes design fixation, ultimately limiting the conceptual space explored by the designer. Similarly, in their study on creative writing, Doshi and Hauser (2024) highlighted that while access to a Large-Language Model (LLM) improved the quality of writing for novices, it resulted in many homogenised outputs, reducing the overall diversity of the stories. McGuire et al. (2024) expands on this and suggests that these limitations occur when the human becomes an editor rather than an active co-creator, highlighting the importance of humans actively directing the creative process. McGuire et al. (2024) further argues that this active involvement can “nurture intrinsic motivation”, a vital component in human creativity Amabile (1985). Our earlier work supports this Krol et al. (2025): In a co-design study with musicians, we found that participants valued owning the creative process and were unwilling to relinquish this control to a machine. Understanding how to design AI-CSTs that promote active involvement in the creative process, by supporting, not replacing, creative skills, is therefore an important avenue of study.
@@
In this work, we explore the themes of ownership and active co-creation in musical AI through a four-week ecological evaluation of a variation system, Rhapsody Refiner Krol et al. (2025). Unlike other symbolic systems Ens and Pasquier (2020); Huang et al. (2019), Rhapsody Refiner does not generate complete new ideas, instead, it relies on the musician to first provide a musical phrase, which it then explores. This forces the musician to be active in both the initial ideation and final creation of the song. The findings from this study demonstrated the tool’s usefulness was not in the generation of complete ideas, but instead moments that sparked inspiration within the composer. Participants highlighted that this made them feel ownership of both the creative process and final composition, noting that this was important to their identities as composers.
@@ 2 Rhapsody Refiner
Rhapsody Refiner is a deep learning-based, music variation system that was co-designed with practising musicians Krol et al. (2025). The tool was inspired by the MidiFormers project [^1] https://github.com/tripathiarpan20/midiformers and utilises MusicBert Zeng et al. (2021), a bidirectional transformer model trained for symbolic music understanding, to perform masked prediction on selected notes. The system supports ideation by generating variations from user-provided phrases and includes functionalities for guiding the output. To generate a variation, users upload a MIDI file and set their desired parameters. The system uses these parameters to guide the masking of random sections of the song, which are then predicted using MusicBert. Participants can choose which aspects of a song to vary - for instance, altering the pitch of notes in a melody or adjusting the dynamics of the entire performance - and can generate multiple variations from the same input. Rhapsody Refiner was packaged as executable software that participants could run without any dependencies.
At the time of evaluation, Rhapsody Refiner had been significantly improved from its previous version and was well suited for studying AI-assisted variation in composition, with co-design workshops providing early evidence of its usability Krol et al. (2025). It is worth noting that for this system to be a complete tool for musicians, it would need to be implemented as a plugin for Digital Audio Workstations (DAW), however this was beyond the scope of this study. The interface for the system can be seen in figure 1 and code for the backend is available on github [^2] https://github.com/SensiLab/MusicVariationBert.
@@ 3 Ecological Evaluation
In order to understand how music variation could fit within composition, we conducted a four week ecological evaluation. Unlike a lab-based evaluation, ecological evaluations allow participants to experiment with software in their own creative environments for an extended period of time. This allows participants to develop a good understanding of the tool they are evaluating and has been used in other music-based studies Habibi et al. (2022); Deruty et al. (2022); Krol et al. (2025).
In total, 8 practising musicians took part in the ecological evaluation. We defined a practising musician as a person that actively engages in the making of music, either professionally or as a hobbyist, and is distinguished from novice musicians who rarely engage in music making Krol et al. (2025). Participants were recruited from known networks in the lab and were provided with a \$50 AUD voucher for their time. This study was also approved by the Monash University Human Research Ethics Committee (MUHREC).
@@
After completing the study, participants took part in a semi-structured interview that explored their usage of the tool. These interviews were transcribed and analysed using an inductive thematic analysis Braun and Clarke (2006) which involved two researchers independently coding the data and coming together to discuss the common themes present in the transcript.
@@ 5.1 Importance of Creative Ownership
Throughout this study, it was clear that ownership, of both the process and artefact, was important to the musical identities of our participants. This concept has been supported in other studies Tchemeube et al. (2023) and this ecological evaluation further demonstrates the importance of designing tools that support this ownership. Because Rhapsody Refiner relies on the user’s creativity to be effective, participants often attributed ownership of the resulting song to themselves, emphasising the effort they invested in shaping the track. Participants noted that they would “genuinely feel bad trying to pass [AI work] as [their] own” (P4) highlighting that a strength of this system is “that it doesn’t totally take charge” (P6). Participants also highlighted a reluctance to adopt tools that would create for them, noting they wanted to feel as though they were the main creative force behind the song.
@@ 5.2 Imperfect Systems Encourage Active Co-Creation
Our four-week ecological evaluation also demonstrated how a system like Rhapsody Refiner, which rarely produces perfect outputs, encouraged the user to be active throughout the co-creative process. The production of moments rather than complete variations meant that participants had to use their own musical expertise to bring these ideas to life. In many cases, these ideas were not completely captured by the system but instead served as a stepping stone for a greater idea “I feel like part of the process is something that happens outside of the variations” (P5). The evaluation also highlighted the value of this interaction model, with participants emphasising the importance of randomness in their creative workflows. This stands in contrast to the prevailing paradigm in AI development, which prioritises the removal of imperfections in pursuit of consistent, ‘clean’ outputs Ouyang et al. (2022).
Naturally, this reliance on the musician has its limitations. Some participants noted that if you have no musical skill or are a novice, it would be harder to extract use from Rhapsody Refiner. While some studies suggest that novices often gain the greatest benefit from AI Doshi and Hauser (2024), the needs of novice and practising musicians are often different Louie et al. (2020); Becker et al. (2024); Krol et al. (2025) and thus require different design choices. Our study suggests that some practising musicians are more willing to embrace imperfect outputs if they lead to better ideas, drawing on their skills to refine and implement those ideas independently. This is in contrast to novice musicians who require more predictability Louie et al. (2020) and who have less ability to turn moments into finished songs.
@@ 5.3 Maintaining Humanity in Music Composition
The making and consuming of music is a unique human practice. It plays an important role in our society as a means for self-expression DeNora (1999), social interaction Rabinowitch and Gill (2021) and therapy Lee (2016). Musicians take pride in making music and despite what some leaders in the technology sector might say [^3] SunoAI CEO on a recent podcast - ”I think the majority of people don’t enjoy the majority of time they spend making music” - enjoy the process of making music Hill (2016). When asked about how P6 felt regarding the emergence of AI in their practice, they responded “Look, I don’t like it from a selfish point of view and also human egocentric point of view, I’d like to think that as a species we’ve got something that’s a little bit more unique. This really challenges that sense of what does it mean to be human?”. This response reflects not only a resistance to adopt AI in music composition, but also deeper anxieties regarding their identity. Understandably, many would hesitate to embrace a technology that threatens to render their own role obsolete, creative or not. However, these anxieties go beyond self preservation and instead question why we would want to automate something that is so innately human.

### v2

@@ 1 Introduction
The music industry is currently experiencing an AI-driven transformation, with various companies now providing complete song generation through prompts (Nayar, 2025). While the revolutionary potential of these technologies remains unclear, a more promising direction has emerged: using AI to support rather than replace human creativity (O’Toole and Ágnes Horvát, 2024; Krol et al., 2025). These supportive applications, known as Artificial Intelligence Creativity Support Tools (AI-CSTs), assist in the creative process, offering varying degrees of agency and control. Some tools are designed to automate routine or technically challenging tasks such as audio-to-midi transcription (Hawthorne et al., 2021) and stem splitting (Hennequin et al., 2020). These systems support the creative process but rarely provide creative ideas. Other tools instead offer more creative involvement, assisting in ideation and even the extension of musical ideas. Notable systems include the Multi-track Music Machine (MMM) (Ens and Pasquier, 2020) which generates controllable multi-track MIDI; Music Transformer (Huang et al., 2019), offering melody harmonisation; and the Anticipatory Transformer (Thickstun et al., 2023), which supports infilling of musical phrases.
@@
Although some research has demonstrated the potential for AI-CSTs to support users (Louie et al., 2020), recent work has highlighted the potential pitfalls of using AI for ideation. For example, Wadinambiarachchi et al. (2024) demonstrated how AI promotes design fixation, ultimately limiting the conceptual space explored by the designer. Similarly, in their study on creative writing, (Doshi and Hauser, 2024) highlighted that while access to a Large-Language Model (LLM) improved the quality of writing for novices, it resulted in many homogenised outputs, reducing the overall diversity of the stories. McGuire et al. (2024) expands on this and suggests that these limitations occur when the human becomes an editor rather than an active co-creator, highlighting the importance of humans actively directing the creative process. (McGuire et al., 2024) further argues that this active involvement can “nurture intrinsic motivation”, a vital component in human creativity (Amabile, 1985). Our earlier work supports this Krol et al. (2025): In a co-design study with musicians, we found that participants valued owning the creative process and were unwilling to relinquish this control to a machine. Understanding how to design AI-CSTs that promote active involvement in the creative process, by supporting, not replacing, creative skills, is therefore an important avenue of study.
@@
In this work, we explore the themes of ownership and active co-creation in musical AI through a four-week ecological evaluation of a variation system, Rhapsody Refiner (Krol et al., 2025). Unlike other symbolic systems (Ens and Pasquier, 2020; Huang et al., 2019), Rhapsody Refiner does not generate complete new ideas, instead, it relies on the musician to first provide a musical phrase, which it then explores. This forces the musician to be active in both the initial ideation and final creation of the song. The findings from this study demonstrated the tool’s usefulness was not in the generation of complete ideas, but instead moments that sparked inspiration within the composer. Participants highlighted that this made them feel ownership of both the creative process and final composition, noting that this was important to their identities as composers.
@@ 2 Rhapsody Refiner
Rhapsody Refiner is a deep learning-based, music variation system that was co-designed with practising musicians (Krol et al., 2025). The tool was inspired by the MidiFormers project [^1] https://github.com/tripathiarpan20/midiformers and utilises MusicBert (Zeng et al., 2021), a bidirectional transformer model trained for symbolic music understanding, to perform masked prediction on selected notes. The system supports ideation by generating variations from user-provided phrases and includes functionalities for guiding the output. To generate a variation, users upload a MIDI file and set their desired parameters. The system uses these parameters to guide the masking of random sections of the song, which are then predicted using MusicBert. Participants can choose which aspects of a song to vary - for instance, altering the pitch of notes in a melody or adjusting the dynamics of the entire performance - and can generate multiple variations from the same input. Rhapsody Refiner was packaged as executable software that participants could run without any dependencies.
At the time of evaluation, Rhapsody Refiner had been significantly improved from its previous version and was well suited for studying AI-assisted variation in composition, with co-design workshops providing early evidence of its usability (Krol et al., 2025). It is worth noting that for this system to be a complete tool for musicians, it would need to be implemented as a plugin for Digital Audio Workstations (DAW), however this was beyond the scope of this study. The interface for the system can be seen in figure 1 and code for the backend is available on github [^2] https://github.com/SensiLab/MusicVariationBert.
@@ 3 Ecological Evaluation
In order to understand how music variation could fit within composition, we conducted a four week ecological evaluation. Unlike a lab-based evaluation, ecological evaluations allow participants to experiment with software in their own creative environments for an extended period of time. This allows participants to develop a good understanding of the tool they are evaluating and has been used in other music-based studies (Habibi et al., 2022; Deruty et al., 2022; Krol et al., 2025).
In total, 8 practising musicians took part in the ecological evaluation. We defined a practising musician as a person that actively engages in the making of music, either professionally or as a hobbyist, and is distinguished from novice musicians who rarely engage in music making (Krol et al., 2025). Participants were recruited from known networks in the lab and were provided with a \$50 AUD voucher for their time. This study was also approved by the Monash University Human Research Ethics Committee (MUHREC).
@@
After completing the study, participants took part in a semi-structured interview that explored their usage of the tool. These interviews were transcribed and analysed using an inductive thematic analysis (Braun and Clarke, 2006) which involved two researchers independently coding the data and coming together to discuss the common themes present in the transcript.
@@ 5.1 Importance of Creative Ownership
Throughout this study, it was clear that ownership, of both the process and artefact, was important to the musical identities of our participants. This concept has been supported in other studies (Tchemeube et al., 2023) and this ecological evaluation further demonstrates the importance of designing tools that support this ownership. Because Rhapsody Refiner relies on the user’s creativity to be effective, participants often attributed ownership of the resulting song to themselves, emphasising the effort they invested in shaping the track. Participants noted that they would “genuinely feel bad trying to pass [AI work] as [their] own” (P4) highlighting that a strength of this system is “that it doesn’t totally take charge” (P6). Participants also highlighted a reluctance to adopt tools that would create for them, noting they wanted to feel as though they were the main creative force behind the song.
@@ 5.2 Imperfect Systems Encourage Active Co-Creation
Our four-week ecological evaluation also demonstrated how a system like Rhapsody Refiner, which rarely produces perfect outputs, encouraged the user to be active throughout the co-creative process. The production of moments rather than complete variations meant that participants had to use their own musical expertise to bring these ideas to life. In many cases, these ideas were not completely captured by the system but instead served as a stepping stone for a greater idea “I feel like part of the process is something that happens outside of the variations” (P5). The evaluation also highlighted the value of this interaction model, with participants emphasising the importance of randomness in their creative workflows. This stands in contrast to the prevailing paradigm in AI development, which prioritises the removal of imperfections in pursuit of consistent, ‘clean’ outputs (Ouyang et al., 2022).
Naturally, this reliance on the musician has its limitations. Some participants noted that if you have no musical skill or are a novice, it would be harder to extract use from Rhapsody Refiner. While some studies suggest that novices often gain the greatest benefit from AI (Doshi and Hauser, 2024), the needs of novice and practising musicians are often different (Louie et al., 2020; Becker et al., 2024; Krol et al., 2025) and thus require different design choices. Our study suggests that some practising musicians are more willing to embrace imperfect outputs if they lead to better ideas, drawing on their skills to refine and implement those ideas independently. This is in contrast to novice musicians who require more predictability (Louie et al., 2020) and who have less ability to turn moments into finished songs.
@@ 5.3 Maintaining Humanity in Music Composition
The making and consuming of music is a unique human practice. It plays an important role in our society as a means for self-expression (DeNora, 1999), social interaction (Rabinowitch and Gill, 2021) and therapy (Lee, 2016). Musicians take pride in making music and despite what some leaders in the technology sector might say [^3] SunoAI CEO on a recent podcast - ”I think the majority of people don’t enjoy the majority of time they spend making music” - enjoy the process of making music (Hill, 2016). When asked about how P6 felt regarding the emergence of AI in their practice, they responded “Look, I don’t like it from a selfish point of view and also human egocentric point of view, I’d like to think that as a species we’ve got something that’s a little bit more unique. This really challenges that sense of what does it mean to be human?”. This response reflects not only a resistance to adopt AI in music composition, but also deeper anxieties regarding their identity. Understandably, many would hesitate to embrace a technology that threatens to render their own role obsolete, creative or not. However, these anxieties go beyond self preservation and instead question why we would want to automate something that is so innately human.